40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Mar 27 23:32:22 2012, frank, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat IMG_0681.jpg
    Reply  Thu Mar 29 00:46:33 2012, frank, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat IMG_0687.JPGIMG_0688.JPG
       Reply  Fri Mar 30 00:50:43 2012, frank, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat compare.png
          Reply  Tue Apr 3 20:56:10 2012, frank, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat beat_2012_04_03.png
             Reply  Wed Apr 4 18:57:11 2012, rana, DailyProgress, Seismic, Seismic coupling from curtain frame to beat 
                Reply  Sun Apr 8 23:35:05 2012, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat beat_ciplot.pngbeat_ciplot.png
                   Reply  Mon Apr 9 23:05:44 2012, tara, DailyProgress, Seismic, Seismic coupling from curtain frame to beat brownian_mult.m.zip
Message ID: 896     Entry time: Tue Apr 3 20:56:10 2012     In reply to: 894     Reply to this: 897
Author: frank, tara 
Type: DailyProgress 
Category: Seismic 
Subject: Seismic coupling from curtain frame to beat 

We measured the beat signal after we fixed the wiring problem. We are close to what we had before, but we still have a scattering light problem.

      Just after we fixed the cable wiring between the table and the frame on 3/30, we measured the beat signal (beat 03/30, blue). There was a hughe scattering noise bump from 100-200 Hz, and it was not stable. We did not see this before (beat 03/12, pink).  We found that there were plenty of dust on the periscope mirrors, lenses in front of RFPD. These optics are in the sensing path, and are more sensitive to scattering light. After cleaning them, the beat signal is better now (beat 04/03). The bump at low frequency is still higher than before, but now it is stable. 

Note that the acoustic peaks around 1kHz are not reduced from fixing the cable, but the bump at low frequency is slightly lower. This is reasonable, the cables should be very soft at high frequency, and do not transfer any vibration from the frame to the table at acoustic bandwidth.

beat_2012_04_03.png

     Other changes we did:

  • replaced the PBS before ACAV AOM with a new one. The old one has defect coating, which we can see by eys, and causes extra back reflection.
  • replaced the PBS in front of RCAV. This one also has high back reflection.
  • re-adjusted the AOM angle. It was tilted down a bit, and this causes extra beam distortion. The beam ellipticity changes from ~0.7 to 0.88 when the AOM is almost parallel to the table surface.
  • replaced translational stage for the curve mirror behind ACAV AOM. Now it should be more stable ( but I don't see any obvious suppression in acoustic peaks yet).
  • sealed the gap between the curtain frame and the plastic covers on top of the table where dust might fall off.
  • change the air spring setup ( I'm not sure about the details, need to ask Frank)

      It is very likely that the when we removed all the cables on the table, dust was stirred up and stuck on the optics, because they are dirtier than usual. We might need to clean more optics on the table. However, from today measurement, it seems that the bump at low frequency might be the effect from the seismic as well (from the feature of the peaks @35 and 55 Hz). Noise budget from seismic will be revisited, also the flat noise at high frequency will be investigated next.

ELOG V3.1.3-